A shopkeeper has 4 types of cooking oil which are Rs.70,Rs.90, Rs100 and 150 per each respectively if he mixed the first two types of oil and last two type in the equal quantities. what is the ratio of oil in the mixture so that the price of the mixture of oil per liter is Rs.98?

Answer by palanisamy(496)

A shopkeeper has 4 types of cooking oil which are Rs.70,Rs.90, Rs100 and 150 per each respectively
Let the shopkeeper mix equal quantities a,a of the first two types of oil.
Let the shopkeeper mix equal quantities b,b of the last two types of oil.
The price of the final mixture of oil per liter is Rs.98
Therefore, we get,
70a+90a+100b+150b = (2a+2b)x 98
160a+250b = 196a+196b
250b-196b = 196a-160a
54b = 36a
b/a = 36/54
b/a = 2/3
a/b = 3/2
The ratio a:b= 3:2